OpenStudy (anonymous):

((sin^2y)/(1-cosy)) I have to rewrite this equation so it is not in a fraction form

OpenStudy (cwrw238):

Hint'; 1/a = a^(-1)

OpenStudy (anonymous):

Try the Pythagorean identity, which states that \[ \sin^2 y + \cos^2 y = 1. \]Using that, you get \[ \frac {\sin^2 y}{1 - \cos y} = \frac {1 - \cos^2 y}{1 - \cos y}. \]Now, recall difference of squares, which states that \[ \frac {1-a^2}{1 - a} = 1 + a. \]Can you finish from here?

OpenStudy (anonymous):

yeah thank you

OpenStudy (anonymous):

You're welcome! A good way of showing your appreciation is by using the Best Response button.

OpenStudy (anonymous):

@sammyplll Thanks!